A chemist mixes 20% HCl with 40% HCl to make 15 liters of solution that is 25% HCl. How many liters of each does he need?

Answer by Alan3354(69443)

t = amt of 20%
f = amt of 40%
-----
t + f = 15 (total solution)
20t + 40f = 25*15 (total acid)